# BiTrajDiff: Bidirectional Trajectory Generation with Diffusion Models for Offline Reinforcement Learning

### v1

@@ 3 Preliminary
Reinforcement Learning (RL) paradigm [42] is formally defined as a Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma,\rho_{0}\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the environment dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\rho_{0}$ denotes the initial state distribution. At each timestep $t$, an agent executes an action $a_{t}$ according to its policy $\pi(\cdot|s_{t})$ in the current state $s_{t}$, transitioning to the subsequent state $s_{t+1}$ with reward $r_{t}$ based on $P$. The objective of the agent is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return: $\pi^{*}=\arg\max_{\pi}\mathcal{J}(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. In contrast to online RL, offline RL [25] constrains the agent from interaction with the environment, relying solely on a static trajectories dataset $\mathcal{D}=\{\tau^{i}\}_{i=1}^{N}$, where $N$ is the number of trajectories and each trajectory $\tau^{i}=\{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\}_{t=0}^{T-1}$ comprises a sequence of state-action-reward tuples.
@@ 4.1 Bidirectional Diffusion Training
where the scalar $\omega\in[0,1]$ denotes the guidance weight of CFG. Setting $\omega$ to 00 degrades the generation to unconditional generation. Conversely, an increased value of $\omega$ promotes a stronger adherence of the generated samples to the provided condition $\mathbf{y}(\tau)$. To train such diffusion denoise models with CFG, the diffusion denoisers are trained with an objective that enables it to predict noise conditioned on $\mathbf{y}(\tau)$ and also unconditionally. This is typically achieved by randomly dropping the condition $\mathbf{y}(\tau)$ during training with probability $p$. Specifically, we have the following loss:
@@ 4.2.1 Bidirectional State Trajectory Generation
$\hat{\tau}_{s}(s_{t})$ =$=\left\{\mathbf{x}^{b}_{0}(\tau)[0:H-1],s_{t},\mathbf{x}^{f}_{0}(\tau)[1:H]\right\}$ (5)
=$=\left\{\hat{s}_{t-H+1},\cdots,\hat{s}_{t-1},s_{t},\hat{s}_{t+1},\cdots,\hat{s}_{t+H-1}\right\},$ (6)
where $H$ is the horizon, and $\hat{s}_{i}$ is the diffusion generated states at MDP timestep $i$. By conditioning on and concatenating at the intermediate anchor state $s_{t}$, we construct a state trajectory, $\hat{\tau}_{s}(s_{t})$, that connects $\hat{s}_{t-H+1}$ and $\hat{s}_{t+H-1}$. This process enables the connection of states that may be disconnected in the original dataset, thereby substantially enriching the diversity of observed behavioral patterns.
@@ 4.2.3 Bidirectional Trajectory Filter
We model the OOD trajectory filter as a novelty detection problem [35], which addresses the identification of data instances that exhibit a significant deviation from the offline training dataset. And we utilize the Isolation Forest model [28] for OOD transition detection. Isolation forest assigns an anomaly score to each data instance, reflecting its susceptibility to isolation with shorter path lengths in randomly partitioned trees. Specifically, we construct the isolation forest on the original offline dataset $\mathcal{D}$, which is able to give each observation $\hat{s}_{i}$ an anomaly score $d(\hat{s}_{i})$. We defined the OOD score of a generated trajectory $d(\hat{\tau})=\sum_{i=t-H+1}^{t+H-2}d(\hat{s}_{i})$. Then we sort $\mathcal{\hat{D}}$ by $d(\hat{\tau})$ and retain the top-$C_{\text{ood}}$ trajectories with the smallest values, where $C_{\text{ood}}$ is a hyperparamter.

### v2

@@ 3 Preliminary
Reinforcement Learning (RL) paradigm [42] is formally defined as a Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma,\rho_{0}\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the environment dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\rho_{0}$ denotes the initial state distribution. At each timestep $t$, an agent executes an action $a_{t}$ according to its policy $\pi(\cdot|s_{t})$ in the current state $s_{t}$, transitioning to the subsequent state $s_{t+1}$ with reward $r_{t}$ based on $P$. The objective of the agent is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return: $\pi^{*}=\arg\max_{\pi}\mathcal{J}(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. In contrast to online RL, offline RL [25] constrains the agent from interaction with the environment, relying solely on a static trajectories dataset $\mathcal{D}=\{\tau^{i}\}_{i=1}^{N}$, where NN is the number of trajectories and each trajectory $\tau^{i}=\{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\}_{t=0}^{T-1}$ comprises a sequence of state-action-reward tuples.
@@ 4.1 Bidirectional Diffusion Training
where the scalar $\omega\in[0,1]$ denotes the guidance weight of CFG. Setting $\omega$ to 0 degrades the generation to unconditional generation. Conversely, an increased value of $\omega$ promotes a stronger adherence of the generated samples to the provided condition $\mathbf{y}(\tau)$. To train such diffusion denoise models with CFG, the diffusion denoisers are trained with an objective that enables it to predict noise conditioned on $\mathbf{y}(\tau)$ and also unconditionally. This is typically achieved by randomly dropping the condition $\mathbf{y}(\tau)$ during training with probability pp. Specifically, we have the following loss:
@@ 4.2.1 Bidirectional State Trajectory Generation
$\hat{\tau}_{s}(s_{t})$ $=\left\{\mathbf{x}^{b}_{0}(\tau)[0:H-1],s_{t},\mathbf{x}^{f}_{0}(\tau)[1:H]\right\}$ (5)
$=\left\{\hat{s}_{t-H+1},\cdots,\hat{s}_{t-1},s_{t},\hat{s}_{t+1},\cdots,\hat{s}_{t+H-1}\right\},$ (6)
where $H$ is the horizon, and $\hat{s}_{i}$ is the diffusion generated states at MDP timestep ii. By conditioning on and concatenating at the intermediate anchor state $s_{t}$, we construct a state trajectory, $\hat{\tau}_{s}(s_{t})$, that connects $\hat{s}_{t-H+1}$ and $\hat{s}_{t+H-1}$. This process enables the connection of states that may be disconnected in the original dataset, thereby substantially enriching the diversity of observed behavioral patterns.
@@ 4.2.3 Bidirectional Trajectory Filter
[…]